OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a yellow marble, do not replace it, and pick another yellow marble?

OpenStudy (anonymous):

2/15 1/15 1/9 3/10

OpenStudy (anonymous):

7

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

7 what

OpenStudy (anonymous):

(total yellow marbles) / (total marbles) gets us the probability each time there are 3 yellow to start with out of 10 , so first we have 3/10 after this we have 2 yellows left and 9 marbles in total so this time its 2/9 for both events to happen we have to multiply them: \[\frac{3}{10} \times \frac{2}{9} = \frac{1}{15}\]

hartnn (hartnn):

probability of picking yellow(3) will be 3/10 since u are not replacing it ,now there are 9 marbles,to get yellow(2),the probability is 2/9 so,overall probab.=3/10*2/9=1/15

OpenStudy (anonymous):

if you still dont understand i could draw a tree-diagram to show you :)

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

i dont know what she means by 7 either...